Hank, a calendar-year taxpayer, uses the cash method of accounting for his sole proprietorship. In late December, he performed $35,000 of legal services for a client. Hank typically requires his clients to pay his bills immediately upon receipt. Assume his marginal tax rate is 32 percent this year and will be 37 percent next year, and that he can earn an after-tax rate of return of 9 percent on his investments.

a. What is the after-tax income if Hank sends his client the bill in December?

b. What is the after-tax income if Hank sends his client the bill in January? Use Exhibit 3.1. (Round your answer to the nearest whole dollar amount.)

EXHIBIT 3-1 Present Value of a Single Payment at Various Annual Rates of Return
4% 5% 6% 7% 8% 9% 10% 11% 12%
Year 1 .962 .952 .943 .935 .926 .917 .909 .901 .893
Year 2 .925 .907 .890 .873 .857 .842 .826 .812 .797
Year 3 .889 .864 .840 .816 .794 .772 .751 .731 .712
Year 4 .855 .823 .792 .763 .735 .708 .683 .659 .636
Year 5 .822 .784 .747 .713 .681 .650 .621 .593 .567
Year 6 .790 .746 .705 .666 .630 .596 .564 .535 .507
Year 7 .760 .711 .665 .623 .583 .547 .513 .482 .452
Year 8 .731 .677 .627 .582 .540 .502 .467 .434 .404
Year 9 .703 .645 .592 .544 .500 .460 .424 .391 .361
Year 10 .676 .614 .558 .508 .463 .422 .386 .352 .322
Year 11 .650 .585 .527 .475 .429 .388 .350 .317 .287
Year 12 .625 .557 .497 .444 .397 .356 .319 .286 .257
Year 13 .601 .530 .469 .415 .368 .326 .290 .258 .229
Year 14 .577 .505 .442 .388 .340 .299 .263 .232 .205
Year 15 .555 .481 .417 .362 .315 .275 .239

c. Based on requirement a and b, should Hank send his client the bill in December or January?
Answer #1

a. What is the after-tax income if Hank sends his client the bill in December?

After tax income = Pre tax income 35000 - Tax ( 35000 * 32 % ) = 35000 - 11200 = $ 23800

b. What is the after-tax income if Hank sends his client the bill in January?

Tax ( 35000 * 37 % ) = 12950

Present value of tax = 12950 * PVIF ( 9 % , 1 year ) = 12950 * 0.917 = 11875

After tax income = Pre tax income 35000 - Tax ( 11875 ) = $ 23125

c. Based on requirement a and b, should Hank send his client the bill in December or January?

In december as proceeds are more in december.